a sttel wire of diameter 1mm can supporta tension of 0.2kn. a cable to support a tension of 20kn should have a diameterv of what order of magnitude
1
Expert's answer
2020-02-07T10:31:35-0500
"\frac{F_2}{F_1}=\frac{A_2}{A_1}=\frac{d_2^2}{d_1^2}"

"\frac{20}{0.2}=\frac{d_2^2}{1^2}"

"d_2=10\ mm"